There are 20 questions in a multiple-choice test. Each question has five choices and one correct answer out of these five choices. A student did not study. He/she answers the questions at random. What are the probabilities that he/she makes (i) more than 10-answer right? (ii) exactly eight right?(iii) at most 7 right?
Answer by stanbon(75887)

There are 20 questions in a multiple-choice test. Each question has five choices and one correct answer out of these five choices. A student did not study. He/she answers the questions at random.
--------------------------------------------
These are binomial problems with n=20, p = 1/2 and q = 1/2
--------------------------------------------
What are the probabilities that he/she makes
(i) more than 10-answers right?
P(11 < x <= 20) = 1 - P(0<= x <= 10) = 1 - 0.588 = 0.412
------------------------------------------
(ii) exactly eight right?
P(x = 8) = 20C8*(1/2)^8*(1/2)^12 = 125970*(1/2)^20 = 0.12
-------------------
(iii) at most 7 right?
P(0 <= x <= 7) = 0.132
==============================
Cheers,
Stan H.